40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Mar 21 09:53:46 2013, Annalisa, Update, Locking, Finding the beat note 
    Reply  Thu Mar 21 10:19:28 2013, Koji, Update, Locking, Finding the beat note 
       Reply  Thu Mar 21 13:11:42 2013, Annalisa, Update, Locking, Finding the beat note Beat_note_setup.jpg
          Reply  Thu Mar 21 13:50:08 2013, Jamie, Update, Locking, incident angles 
             Reply  Thu Mar 21 14:02:45 2013, Annalisa, Update, Locking, incident angles 
                Reply  Fri Mar 22 23:23:38 2013, Annalisa, Update, Locking, Beat note still missing Beat_note_-_new_setup.jpg
Message ID: 8333     Entry time: Fri Mar 22 23:23:38 2013     In reply to: 8329
Author: Annalisa 
Type: Update 
Category: Locking 
Subject: Beat note still missing 

[Annalisa, Manasa, Koji]

I updated the setup for the beat note. The main reason is that I needed to keep the ADJ to 0 in way to operate at the nominal laser power.

Now the input power of the laser is increased (about 315 mW) and needs to be dumped so as not to exceed the PD threshold of 1mW. 

Moreover, a lens has been added to match the two beams size.

A BS has been removed from the PSL pick-off beam path, so the PSL power hitting the BS is now about 100 uW, and the total power on the PD is 0.7mW.

I also verified that both the beams are S polarized.

To find the beat note, the laser temperature has been varied  through the laser controller and not adding a Voltage with the power supply.

A range of temperature of 30 degC has been spanned, but we suppose there should be some calibration problem with the controller, since set temperature is not the same as Laser temperature on the display. 

Anyway, no beat note has been found up to now.


An external monitor has to be added to check the real temperature of the crystal.

The next possible plan is to vary the PSL temperature and try to find the beat note. 


P.S.: The HEWELETT PACKARD 8591E spectrum analyzer works! The monitor only took some time to turn on!


Attachment 1: Beat_note_-_new_setup.jpg  145 kB  | Hide | Hide all
ELOG V3.1.3-